OpenStudy (anonymous):

Solve the equation for a. 55a – 11 = 44a

mathslover (mathslover):

Add 11 both sides.

mathslover (mathslover):

\(55a - 11 +\color{blue}{11} = 44a + \color{blue}{11} \) What do you get ?

OpenStudy (anonymous):

66a=55a

OpenStudy (uri):

44a -55a =11 11a =11 a =11/11 a =1 ans :D